2x^2-x=-5 solution set= ?
Answer by edjones(8007)

2x^2-x=-5
2x^2-x+5=0
x^2-x/2+5/2=0 divide both sides by 2 so the coefficient of x^2 is 1.
x^2-x/2 =-5/2
x^2 - x/2 + 1/16 = -5/2 + 1/16 Complete the square.
(x - 1/4)^2 = -39/16
x - 1/4 = +- sqrt(39)/4 * i Take the sqrt of each side.
x=1/4 + sqrt(39)/4 * i, x=1/4 - sqrt(39)/4 * i
.
Ed